OpenStudy (anonymous):

The peak intensity of radiation from a star named Sigma is 2 x 106 nm. What is the average surface temperature of Sigma rounded to the nearest whole number? 1.45 K 58 K 1,450 K 5.8 x 106K

OpenStudy (anonymous):

Did you mean 2E6? 2\(\cdot 10^6\)?

OpenStudy (anonymous):

\[λ_{\max}=\frac{b}{T}\] You can do this without having to convert the constant "b" over from meters\(\cdot\)Kelvin, into nanometer\(\cdot\)electron-Volts by converting your wavelength to meters first, so the units cancel out. b, for the max of Planck's radiation curve here = 0.00289777 m\(\cdot\)K Can you find the answer from here? (hint: it has to be very cold for a wavelength that large)